### v1

# Predicting function of evolutionarily implausible DNA sequences

Abstract
Genomic language models (gLMs) show potential for generating novel, functional DNA sequences for synthetic biology, but doing so requires them to learn not just evolutionary plausibility, but also sequence-to-function relationships. We introduce a set of prediction tasks called Nullsettes, which assesses a model’s ability to predict loss-of-function mutations created by translocating key control elements in synthetic expression cassettes. Across 12 state-of-the-art models, we find that mutation effect prediction performance strongly correlates with the predicted likelihood of the nonmutant. Furthermore, the range of likelihood values predictive of strong model performance is highly dependent on sequence length. Our work highlights the importance of considering both sequence likelihood and sequence length when using gLMs for mutation effect prediction.
Machine Learning, ICML
1 Introduction
Genomic language models (gLMs) learn a probability distribution over DNA sequences, representing the evolutionary plausibility of genomic sequences (Benegas et al., 2025b; Consens et al., 2025). For example, gLMs would assign higher likelihood values for regulatory motifs that occur frequently in natural genomic sequences compared to random sequences.
Evolutionary plausibility can be a useful proxy for biological function. In both genomic and protein language models (gLMs and pLMs), higher model likelihoods often correlate with improved functional properties. For example, pLM likelihoods can guide the design of higher affinity antibodies (Hie et al., 2024), more efficient base editors (He et al., 2024), and brighter fluorescent proteins (Zhang et al., 2025). Similarly, gLM likelihoods enable genome-wide prediction of variant effects (Benegas et al., 2023) and 5’ untranslated region (UTR) optimization (Chu et al., 2024).
However, current evaluations for gLMs on mutant effect prediction focus primarily on natural sequences. In contrast, synthetic biology often requires the design of functional sequences with little or no evolutionary precedent, such as sequences that confer novel functions, avoid crosstalk with native machinery, or push expression levels beyond natural limits. Examples include ultra-strong synthetic promoters (Schlabach et al., 2010), miRNA-based regulatory circuits that achieve dosage-compensated gene expression using elements orthogonal to native miRNAs (Du et al., 2024), and engineered metabolic pathways for production of small molecule drugs (Yan et al., 2023). For these applications, it is unclear whether gLMs trained on natural genomes can generalize to synthetic constructs, especially in the absence of deep mutational scanning data for DNA. If successful, gLMs could support more systematic and scalable approaches to genetic design.
Figure 1: Overview of Nullsettes. A) Schematic examples showing transcription (i) and translation (ii) becoming impaired when the order of control element such as terminator (iii) or stop codon (iv) are altered. B) Nullsettes generates 11 and 19 mutant variants for each eukaryotic and prokaryotic expression cassette, respectively, by translocating control elements, and evaluates how well gLMs can predict these loss-of-function mutations based on changes in log-likelihood.
We introduce a benchmark suite, Nullsettes, to evaluate gLMs on their ability to predict loss-of-function (LOF) mutations in synthetic expression cassettes. Nullsettes leverages well-established mechanisms of gene expression to systematically introduce virtual LOF mutations by rearranging key control elements, such as promoters and start codons, within functional cassettes. Using functional cassettes (referred to as nonmutant) curated from Massive Parallel Reporter Assay (MPRA) data  (Lagator et al., 2022; de Boer et al., 2020; Kosuri et al., 2013; Zahm et al., 2024), including those with randomly generated, low-likelihood promoters, we show that Evo-2-7B outperforms 11 other state-of-the-art models in mutant effect prediction. However, all gLMs exhibit a sharp decline in predictive accuracy as the likelihood of the nonmutant sequence decreases. Moreover, the likelihood range at which each model performs optimally varies strongly with sequence length. These results suggest that both likelihood and length must be considered when applying gLMs to guide genetic design.
2 Methods
2.1 Nullsettes Construction
To systematically evaluate gLMs’ ability for zero-shot functional prediction on expression cassettes, we proposed Nullsettes, a suite of virtual mutants that perturb transcription and translation of an expression cassette by shuffling the order of six key control elements. These elements are promoter, ribosome binding site (RBS), start codon, coding sequence (CDS), stop codon, and terminator.
In the canonical configuration of an expression cassette (Figure 1A i-ii), the proper ordering of promoter–start codon–CDS–stop codon–terminator from 5’ to 3’ end enables proper transcription (i) and translation (ii). However, translocations that reposition critical elements, such as placing the terminator between the promoter and the CDS (iii) or the CDS downstream of the stop codon (iv), result in transcriptional or translational failure, respectively.
Nullsettes consists of single-element translocations that effectively eliminates any production of functional mRNA and protein (Figure 1B). Among the set of all possible single-element translocations, we subset for mutants where 1) the CDS expression cannot be rescued by any flanking sequences outside the cassette, 2) expression machinery can still act to generate useless products. These virtual mutant cassettes collectively make up the Nullsettes, which consists of 11 mutations for eukaryotic and 19 mutations for prokaryotic cassettes (see Appendix Table 3 and 4 for full list). The 8 additional mutants for prokaryote cassettes are due to translocating the RBS. Given a functional cassette, we then compare its gLM log-likelihood (LL) with its corresponding Nullsettes to assess how well a gLM can perform zero-shot functional prediction by consistently predicting lower likelihoods for the Nullsettes.
2.2 Functional cassette curation
To assess mutant effect prediction using Nullsettes, we curate expression cassettes from public MPRA datasets  (Lagator et al., 2022; de Boer et al., 2020; Kosuri et al., 2013; Zahm et al., 2024). We specifically use cassettes with strong gene expression but low gLM likelihood representing low evolutionary plausibility. These cassettes have low likelihood due to containing elements from evolutionarily distant organisms. For example, the CDS, green fluorscent protein (GFP), is from the jellyfish Aequorea victoria, whereas promoter and terminators are sourced from bacteria, human and mice. Furthermore, cassettes curated from Lagator dataset and deBoer pTpA dataset use random DNA sequences as promoters, which further lowers gLM likelihood compared to those with natural promoters/motifs (See Appendix A.3). In the paper, we name datasets with random promoters with the suffix (“Low”) and those with more natural promoters with the suffix (“High”) (Figure 2).
Figure 2: Relationship between gLM likelihood and zero-shot performance on Nullsettes prediction. A) Each point represents an expression cassette. Scatterplots show the Nullsettes prediction performance of Evo-2-7B and Nucleotide Tranformer-2.5B-MS for sequences from five datasets. B) Heatmap showing correlation between the gLM log likelihood (LL) of a sequence and gLM success rate on Nullsettes prediction, across five datasets and 12 models. C) Stacked bar plot showing success rate of models computed using all sequences in each dataset, representing the proportion of Nullsettes mutants that a model consistently outputs a lower LL compared to the nonmutant.
2.3 Baseline Models
We benchmarked a diverse set of 12 self-supervised genomic foundation models that represent current state-of-the-art approaches to DNA language modeling. In general, the models can be categorized based on their tokenization schemes (fixed-length k-mers, single-nucleotide tokens, byte-pair encoding, and hybrid schemes), pretraining strategies (masked language modeling (Devlin et al., 2019) vs. autoregressive modeling (Brown et al., 2020)), training corpus diversity (human references, plant genomes, multispecies genomes, and large-scale metagenomic assemblies), and architectural paradigms (CNN-based models, transformer-based models (Consens et al., 2025), and emerging architectures like StripedHyena (Poli et al., 2023; Ku et al., 2025) and Mamba (Gu & Dao, 2023)). Detailed model specifications are concluded in Appendix Table 1 and 2.
2.4 Evaluation Metrics
We computed the mean base-pair log-likelihood (LL) as the sequence-level LL score. To account for the distinct pretraining objectives of causal language models (CLMs) and masked language models (MLMs), we applied different LL computation strategies, as detailed in Appendix B.1. After obtaining the LL distributions for both the original cassette and its virtual mutants (Nullsettes), we performed a one-sided paired permutation test to assess whether the mutant cassettes exhibit a significant decrease in LL compared to the original sequence (Figure 1B). Additional methodological details are provided in Appendix B.2.
3 Results
We benchmarked these gLMs and their variants across five datasets, as shown in Appendix C. Representative models were selected for further analysis in the following sections.
3.1 Zero-shot Nullsettes prediction performance deteriorates for sequences with low gLM likelihood
The evolutionary plausibility of a sequence, as predicted by a gLM, strongly correlates with the gLM’s performance in predicting mutation effects for the sequence. Figure 2A compares gLM prediction performance of individual functional cassettes and their predicted LL for two models, Evo-2-7B and NucleotideTransformer-2.5B-MS. Here success rate for each cassette is defined as the proportion of Nullsettes (11 in eukaryote, 19 in prokaryote) whose LL score is lower than the original, nonmutant cassette. For both models, we observe strong correlations between the nonmutant LL score and the gLM’s ability to identify LOF mutants. For example, for the E.coli (Low) dataset (Lagator et al., 2022), zero-shot Nullsettes prediction with NT-2.5B-MS is nearly impossible when the nonmutant LL is below -25, and easily accomplished for nonmutants with LL above -22. In fact, Figure 2B shows that all models, despite variations in architecture and training, exhibits a positive correlation between sequence likelihood and mutant prediction performance.
The Evo-2-7B model achieves the best overall performance in zero-shot Nullsettes prediction. Evaluating zero-shot prediction using all sequences within a dataset (See Methods), Figure 2C shows that Evo-2-7B achieves the best overall performance across all five datasets. Specifically for the E.coli and yeast MPRA data using random promoters (left), which are less evolutionarily plausible than more natural sequences shown on the right (Appendix Figure 4).
Figure 3: Optimal likelihood range for Nullsettes prediction varies with sequence length. A) Length in base pairs and normalized gLM log likelihood of all nonmutant cassettes, averaged across all 12 models. B) Contour of scatter plots similar to those in Figure 2A, colored by dataset, dotted vertical lines are rough estimates of optimal LL value for sequences from different datasets. Human data for Evo 1.5 is cut off for better visual due to large LL score.
3.2 Optimal likelihood range predictive of strong model performance varies with sequence length
Although mutant prediction performance correlates with sequence likelihood, the range of LL scores for which a model performs well depends highly on the length of the DNA sequence. The range of LL scores for which models achieve strong performance is higher for longer sequences, partly due to the fact that gLMs tend to assign higher LL scores to longer cassettes (Figure 3A). For example, Figure 3B shows that the NT-2.5B-MS model achieves strong performance for cassettes with log likelihood above $-22$ for both E. coli datasets, as indicated by the gray dotted line. However, for sequences from the Human GPRA dataset which are $2.5$-folds longer, only LL score above $-17$ achieved strong performance. Similarly sequences from the yeast (low) dataset are $30\%$ longer compared to sequences from the E. coli datasets. Thus, for GPN-promoter and Evo-1.5, the optimal likelihood range is also significant higher for the yeast (low) dataset.
Altogether, there does not appear to be a single optimal likelihood range within which a model would be able to make zero-shot mutation effect predictions, as sequence properties such as length can have a significant effect on optimal likelihood range.
4 Discussion
In this study, we introduce Nullsettes, the first systematic benchmark for assessing the ability of genomic language models (gLMs) to predict loss-of-function mutations in synthetic expression cassettes. By enabling simple in silico translocation of sequences, Nullsettes facilitates generalizable, cassette-independent evaluations of gLM functional performance. Beyond general performance benchmarking, our results reveal a strong correlation between gLM performance on mutation prediction tasks and the model’s predicted likelihood of the original, non-mutant sequence. Moreover, we demonstrate that the optimal likelihood range for accurate mutant discrimination is not fixed but varies significantly with sequence length. Together, we hope this benchmark and its findings will inspire the development of gLMs that move beyond evolutionary priors to support the rational design of functional, out-of-distribution genetic sequences.
Software and Data
The source code for benchmarking gLM on Nullsettes is publicly available at: https://github.com/cellethology/GLM-Nullsette-Benchmark.
Acknowledgements
The authors thank members of the Cell Ethology Lab for valuable comments and suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of genomic language model and its application in synthetic biology. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Datasets
A.1 Prokaryote
To evaluate the regulatory grammar of prokaryotes on genomic language models, we utilized two benchmark datasets derived from synthetic expression constructs in Escherichia coli. The first, the Kosuri dataset (Kosuri et al., 2013), comprises combinatorial assemblies of promoters and ribosome binding sites (RBS) upstream of a superfolder GFP (sfGFP) reporter, designed to dissect transcriptional and translational contributions to gene expression. The second, the Lagator dataset (Lagator et al., 2022), features a large-scale library of random promoter sequences with experimentally measured activity, enabling evaluation of models’ ability to generalize to highly diverse, out-of-distribution regulatory inputs.
A.2 Eukaryote
To evaluate the capacity of genomic language models to generalize to eukaryotic regulatory logic, we utilized two large-scale massively parallel reporter assay (MPRA) datasets. The Zahm dataset (Zahm et al., 2024) comprises a library of 6,144 synthetic promoters constructed by combining transcriptional response elements (TREs) from 229 human and mouse transcription factors with minimal promoters. These constructs were assayed across multiple human cell lines and stimulatory conditions to quantify dynamic, stimulus-specific gene expression. The deBoer dataset (de Boer et al., 2020) includes a comprehensive collection of 100 million randomized promoter sequences in yeast, enabling high-resolution dissection of cis-regulatory grammar under out-of-distribution scenario.
A.3 Promoter selection
To construct decoupled expression cassettes, we selected 1,500 promoters or promoter–RBS pairs from four distinct datasets.
A.3.1 Kosuri and Lagator datasets
For the Kosuri dataset (Kosuri et al., 2013), we selected candidate promoter–RBS pairs with protein expression levels exceeding $\mu+1.5\sigma$ (10,165) across the full distribution. In contrast, given the limited number of highly active constructs in the Lagator dataset, we directly selected the top 1,500 promoter sequences ranked by protein output. We then compared the log-likelihood (LL) distributions of promoter–RBS groups using genomic language models including Evo1 8k, GENERator 3B, GPN, and Nucleotide Transformer 2.5B multispecies. As the Lagator dataset (Lagator et al., 2022) consists of randomly generated promoters, whereas the Kosuri library is more rationally designed, we prioritized Kosuri promoter–RBS pairs whose LL distributions surpassed those from Lagator (Appendix Figure 4(a)), yielding 1,500 expression cassettes from each dataset.
A.3.2 deBoer dataset
deBoer dataset (de Boer et al., 2020) contains two promoter libraries: Abf1TATA and pTpA. Abf1TATA is designed by embedding conserved transcription factor binding sites such as Abf1 and a canonical TATA box, mimicking features of natural yeast promoters; while pTpA consists of synthetic promoters constructed with a simple poly-T–poly-A architecture, lacking specific transcription factor motifs and serving as a minimal, randomized control library. Similar to Kosuri vs Lagator, we selectively constructed 1,500 active promoters for each dataset based on LL distribution where pTpA is left-shifted comparing to Abf1TATA as Abf1TATA is more conservative (Appendix Figure 4(b)). The source can be found at NCBI’s GEO: GSE104878111https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE104878.
A.3.3 Zahm dataset
The Zahm dataset (Zahm et al., 2024) comprises synthetic promoters built by coupling one of three minimal promoters—minCMV, minProm, or minTK—with diverse transcriptional response element (TRE) units. To construct a representative set of mammalian expression cassettes, we selected the top 1,500 promoter-TRE combinations exhibiting the highest transcriptional activity. The source can be found at NCBI’s GEO: GSE271608222https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE271608.
(a) Kosuri vs Lagator
(b) deBoer TATA vs deBoer pTpA
Figure 4: Comparison of promoter log-likelihood distribution
A.4 Virtual mutant
Translocation mutants were systematically engineered by permuting the canonical order of key regulatory and coding elements—promoter, ribosome binding site (RBS), start codon, coding sequence (CDS), stop codon, and terminator in prokaryotic systems; and promoter, start codon, CDS, stop codon, and terminator in eukaryotic systems. Each permutation aims to disrupt the natural transcription-translation flow, creating syntactically or functionally invalid constructs. A total of 19 translocation variants were designed for prokaryotic systems and 11 for eukaryotic systems. All constructs were rendered completely non-functional by design. All translocation combinations can be found in Appendix Table 3 for prokaryote and Appendix Table 4 for eukaryote.
Appendix B Evaluation Metrics
B.1 Mean Log-likelihood
For casual language models (CLM), the log-likelihood of a nucleic acid sequence $X=(x_{1},x_{2},...,x_{n})$ is computed using logits output from the model. Since CLM operates in an autogressive manner, each token $x_{t}$ is predicted based on all previous token $x<t$. Given the model’s ${logits}_{t}$ at each position $t$, the probability of the ground-truth token is obtained via softmax:
$P(x_{t}\mid x_{<t})=\frac{\exp(\text{logits}_{t,x_{t}})}{\sum_{v\in V}\exp(\text{logits}_{t,v})}$ (1)
where $V$ represents the vocabulary. The mean log-likelihood is then computed as:
$\text{LL}_{\text{CLM}}(X)=\frac{1}{n-1}\sum_{t=2}^{n}\log P(x_{t}\mid x_{<t})$ (2)
Masked language models (MLMs) differ from autoregressive models in their inference behavior, as they are not inherently designed for sequential generation. To address this, Wang & Cho (2019) introduced the pseudo log-likelihood (PLL) approach, wherein each token in a sequence is masked individually to compute token-wise conditional probabilities. More recently, Gordon et al. (2024) proposed Single-Inference Pseudo Log-Likelihood, an efficient approximation that enables linear-time inference for BERT-style MLMs by exploiting training-time masking dynamics.
Rather than computing pseudo log-likelihood (PLL), we assessed relative changes in mean token-wise log-probabilities derived from unmasked logits for all MLM-based gLMs. This method, though lacking strict probabilistic interpretation, provides a scalable and effective proxy for quantifying model sensitivity to syntactic or functional disruptions. The formula is as follows:
Given the ${logits}_{t}$ at each position $t$, the probability of the correct token is computed as:
$P(x_{t}\mid X)=\frac{\exp(\text{logits}_{t,x_{t}})}{\sum_{v\in V}\exp(\text{logits}_{t,v})}$ (3)
Then the log-likelihood for the entire sequence is computed over all positions:
$\text{LL}_{\text{MLM}}(X)=\frac{1}{n}\sum_{t=1}^{n}\log P(x_{t}\mid X)$ (4)
where all positions contribute to the likelihood computation since no masking is performed during inference.
B.2 One-sided Paired Permutation Test
To assess significance between paired conditions (e.g., original vs. mutant), we performed as one-sided paired permutation test. Given differences $d_{i}={x_{i}}^{mutant}-{x_{i}}^{original}$ for each pair $i$, we computed the observed mean $\bar{d}_{\text{obs}}$. Under the null hypothesis of no effect, signs of $d_{i}$ are exchangeable. We generated $N=10,000$ random permutations by sampling $s_{i}\in\{-1,1\}$ and computing $\bar{d}^{(j)}=\frac{1}{n}\sum_{i=1}^{n}s_{i}d_{i}$. The one-sided p-value was calculated as:
$p=\frac{1}{N}\sum_{j=1}^{N}\mathbb{I}\left(\bar{d}^{(j)}\leq\bar{d}_{\text{obs}}\right)$ (5)
for the alternative hypothesis that the mutant is smaller than the original.
Appendix C Model Performance Evaluation
We systematically evaluated the impact of tokenization strategies and model sizes on the performance of gLMs. As shown in Appendix Figure 5, we benchmarked the PDLLM model series (Liu et al., 2025a), which differ only in their tokenization choices. Our analysis revealed that the choice of tokenization profoundly influences model success rates: byte pair encoding (BPE) shows a stable performance across models. In contrast, single-token schemes performed poorly, while the 6-mer representation exhibited intermediate performance, with model-specific variability. To assess the effect of model size, we compared models that vary only in parameter size. Notably, we found no consistent relationship between model size and task performance: in several cases, smaller models performed on par with, or even outperformed, their larger counterparts.
Figure 5: Comparative evaluation of gLMs across tokenization methods (left) and model sizes (right).
Furthermore, while representative models were selected for the main analysis, we conducted an extensive evaluation of various model versions across all datasets. As shown in Appendix Figure 6, the Evo series models (Nguyen et al., 2024; Merchant et al., 2024; Brixi et al., 2025) consistently outperformed all others in the Nullsettes tasks, with nearly every Evo variant ranking at the top. This advantage likely reflects the superior architectural design and scaling strategy employed by the Evo models. Among non-Evo models, METAGENE-1 (Liu et al., 2025b) achieved the highest performance, likely due to its training on over 1.5 trillion base pairs of DNA and RNA sequences derived from wastewater metagenomes. The unique breadth and diversity of metagenomic data may provide this model with an enhanced ability to generalize across sequence contexts. The Nucleotide Transformer (NT) (Dalla-Torre et al., 2024) series followed, ranking third overall. These findings highlight the impact of both model architecture and training data diversity on generalization performance across regulatory sequence prediction tasks.
Figure 6: The number of cumulative successful predictions made by each model series across four datasets: Abf1TATA and pTpA (from deBoer), Zahm, Kosuri, and Lagator.
Model name Pretraining dataset Pretraining method Tokenization Architecture Input length
Nucleotide Transformer 2.5B multispecies (Dalla-Torre et al., 2024) 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e. roughly 29B tokens. MLM 6-mers Transformer, 2.5B params 1000 bp
Nucleotide Transformer 500M multispecies (Dalla-Torre et al., 2024) 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e. roughly 29B tokens. MLM 6-mers Transformer, 500M params 1000 bp
Nucleotide Transformer 500M Human ref (Dalla-Torre et al., 2024) GRCh38 human reference genome, resulting in 3B nucleotides, i.e. roughly 500M 6-mer tokens. MLM 6-mers Transformer, 500M params 1000 bp
Evo1 8k (Nguyen et al., 2024) OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context
Evo1 131k (Nguyen et al., 2024) OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 131,072 context using Evo1 8k as the base model
Evo1.5 (Merchant et al., 2024) 50% increase in training data comparing to Evo1 CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context using Evo1 8k as the base model
Evo2 1b base (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 1B params Pretrained with 8192 context length
Evo2 7b base (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 8192 context length
Evo2 7b (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 1M context using Evo2 7b base as the base model
GENERator 3b (Wu et al., 2025) Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate. CLM 6-mer Llama-based decoder, 3B params Pretrained with a context length of 98k bp
GENERator 1.2b (Wu et al., 2025) Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate. CLM 6-mer Llama-based decoder, 1.2B params Pretrained with a context length of 98k bp
METAGENE-1 (Liu et al., 2025b) 1.5T base pairs of DNA and RNA sequences from human wastewater samples. CLM Byte-pair encoding Llama-2, 7B params Pretrained with 512 sequence length
Table 1: Genomic language model details
Model name Pretraining dataset Pretraining method Tokenization Architecture Input length
Caduceus-ph (Schiff et al., 2024) Human reference genome (35M tokens / nucleotide base pairs). MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (reverse complement augmentation)
Caduceus-ps (Schiff et al., 2024) Human reference genome (35M tokens / nucleotide base pairs). MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (No reverse complement augmentation)
GENA-LM human (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly. MLM Byte-pair encoding BERT/BigBird, 110M params About 4,500 nucleotides (512 BPE tokens)
GENA-LM multi (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly augmented by sampling mutations from 1000-genome SNPs (gnomAD dataset) + Multispecies genomes from ENSEMBL release 108. MLM Byte-pair encoding BERT/BigBird, 110M params About 4,500 nucleotides (512 BPE tokens)
GENA-LM large (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly. MLM Byte-pair encoding BERT/BigBird, 336M params About 4,500 nucleotides (512 BPE tokens)
GPN (Benegas et al., 2023) Brassicales reference genome from NCBI Genome. Took the union of exons (with a small intronic flank), promoters (1,000 bp upstream of transcription start sites) as well as an equivalent amount of random windows from the whole genome. MLM Single-nucleotide tokens CNN 512 bp
GPN-Promoter (Benegas et al., 2025a) Animal promoter, genomes of 434 animal species. MLM Single-nucleotide tokens ByteNet, 152M Pretrained on 512 bp sequences centered at TSSs of protein-coding genes
DNABERT 2 (Zhou et al., 2023) Human genome dataset (2.75B nucleotide bases) + Multispecies genome dataset (from 135 species, spread across 6 cetogories). The dataset includes 32.49B nucleotides bases, excluding all sequence with N and retain only ATCG. MLM Byte-pair encoding BERT, 117M params Pretrained on 700 bp length sequences
gLM2 (Cornman et al., 2024) OMG: encodes genomic scaffolds with both amino-acid (CDS) and DNA tokens (315B tokens). MLM Char-level tokens (DNA: lowercase, AA: uppercase) Transformer, 650M params Pretrained with a 4096 token context window
PDLLM-DNAMamba-6mer (Liu et al., 2025a) Plant reference genomes CLM 6-mer SSM, 130M params Pretrained max token length is 512
Table 2: Genomic language model details (continued)
Mutant ID Description
Translocation-1 CDS - Promoter - RBS - Start codon - Stop codon - Terminator
Translocation-2 Promoter - CDS - RBS - Start codon - Stop codon - Terminator
Translocation-3 Promoter - RBS - CDS - Start codon - Stop codon - Terminator
Translocation-4 Promoter - RBS - CDS - Stop codon - Start codon - Terminator
Translocation-5 Promoter - RBS - CDS - Stop codon - Terminator - Start codon
Translocation-6 Promoter - RBS - Start codon - Stop codon - CDS - Terminator
Translocation-7 Promoter - RBS - Start codon - Stop codon - Terminator - CDS
Translocation-8 Promoter - RBS - Start codon - Terminator - CDS - Stop codon
Translocation-9 Promoter - RBS - Terminator - Start codon - CDS - Stop codon
Translocation-10 Promoter - Start codon - CDS - RBS - Stop codon - Terminator
Translocation-11 Promoter - Start codon - CDS - Stop codon - RBS - Terminator
Translocation-12 Promoter - Start codon - CDS - Stop codon - Terminator - RBS
Translocation-13 Promoter - Start codon - RBS - CDS - Stop codon - Terminator
Translocation-14 Promoter - Terminator - RBS - Start codon - CDS - Stop codon
Translocation-15 RBS - Promoter - Start codon - CDS - Stop codon - Terminator
Translocation-16 RBS - Start codon - CDS - Promoter - Stop codon - Terminator
Translocation-17 RBS - Start codon - CDS - Stop codon - Promoter - Terminator
Translocation-18 RBS - Start codon - Promoter - CDS - Stop codon - Terminator
Translocation-19 Start codon - Promoter - RBS - CDS - Stop codon - Terminator
Table 3: Prokaryotic virtual mutant cases. Translocation-3,4,5 cannot be compensated by a later start codon as there are no other in-frame start codon.
Mutant ID Description
Translocation-1 CDS - Promoter - Start codon - Stop codon - Terminator
Translocation-2 Promoter - CDS - Start codon - Stop codon - Terminator
Translocation-3 Promoter - CDS - Stop codon - Start codon - Terminator
Translocation-4 Promoter - CDS - Stop codon - Terminator - Start codon
Translocation-5 Promoter - Start codon - Stop codon - CDS - Terminator
Translocation-6 Promoter - Start codon - Stop codon - Terminator - CDS
Translocation-7 Promoter - Start codon - Terminator - CDS - Stop codon
Translocation-8 Promoter - Terminator - Start codon - CDS - Stop codon
Translocation-9 Start codon - CDS - Promoter - Stop codon - Terminator
Translocation-10 Start codon - CDS - Stop codon - Promoter - Terminator
Translocation-11 Start codon - Promoter - CDS - Stop codon - Terminator
Table 4: Eukaryotic virtual mutant cases. Translocation-2,3,4 cannot be compensated by a later start codon as there are no other in-frame start codon.

### v2

@@
 Appendix C Model Performance Evaluation
-We systematically evaluated the impact of tokenization strategies and model sizes on the performance of gLMs.
-As shown in Appendix Figure 5, we benchmarked the PDLLM model series (Liu et al., 2025a), which differ only in their tokenization choices.
-Our analysis revealed that the choice of tokenization profoundly influences model success rates: byte pair encoding (BPE) shows a stable performance across models.
-In contrast, single-token schemes performed poorly, while the 6-mer representation exhibited intermediate performance, with model-specific variability.
-To assess the effect of model size, we compared models that vary only in parameter size.
-Notably, we found no consistent relationship between model size and task performance: in several cases, smaller models performed on par with, or even outperformed, their larger counterparts.
-Figure 5: Comparative evaluation of gLMs across tokenization methods (left) and model sizes (right).
-Furthermore, while representative models were selected for the main analysis, we conducted an extensive evaluation of various model versions across all datasets.
-As shown in Appendix Figure 6, the Evo series models (Nguyen et al., 2024; Merchant et al., 2024; Brixi et al., 2025) consistently outperformed all others in the Nullsettes tasks, with nearly every Evo variant ranking at the top.
+While representative models were selected for the main analysis, we conducted an extensive evaluation of various model versions across all datasets.
+As shown in Appendix Figure 5, the Evo series models (Nguyen et al., 2024; Merchant et al., 2024; Brixi et al., 2025) consistently outperformed all others in the Nullsettes tasks, with all the Evo variants ranking at the top.
 This advantage likely reflects the superior architectural design and scaling strategy employed by the Evo models.
@@
 MLM Byte-pair encoding BERT, 117M params Pretrained on 700 bp length sequences
-gLM2 (Cornman et al., 2024) OMG: encodes genomic scaffolds with both amino-acid (CDS) and DNA tokens (315B tokens).
-MLM Char-level tokens (DNA: lowercase, AA: uppercase) Transformer, 650M params Pretrained with a 4096 token context window
 PDLLM-DNAMamba-6mer (Liu et al., 2025a) Plant reference genomes CLM 6-mer SSM, 130M params Pretrained max token length is 512

### v3

@@
-Predicting function of evolutionarily implausible DNA sequences
+Evaluating DNA function understanding in genomic language models using evolutionarily implausible sequences
 Abstract
-Genomic language models (gLMs) show potential for generating novel, functional DNA sequences for synthetic biology, but doing so requires them to learn not just evolutionary plausibility, but also sequence-to-function relationships.
-We introduce a set of prediction tasks called Nullsettes, which assesses a model’s ability to predict loss-of-function mutations created by translocating key control elements in synthetic expression cassettes.
-Across 12 state-of-the-art models, we find that mutation effect prediction performance strongly correlates with the predicted likelihood of the nonmutant.
-Furthermore, the range of likelihood values predictive of strong model performance is highly dependent on sequence length.
-Our work highlights the importance of considering both sequence likelihood and sequence length when using gLMs for mutation effect prediction.
-Machine Learning, ICML
-1 Introduction
-Genomic language models (gLMs) learn a probability distribution over DNA sequences, representing the evolutionary plausibility of genomic sequences (Benegas et al., 2025b; Consens et al., 2025).
-For example, gLMs would assign higher likelihood values for regulatory motifs that occur frequently in natural genomic sequences compared to random sequences.
-Evolutionary plausibility can be a useful proxy for biological function.
+Genomic language models (gLMs) hold promise for generating novel, functional DNA sequences for synthetic biology.
+However, realizing this potential requires models to go beyond evolutionary plausibility and understand how DNA sequence encodes gene expression and regulation.
+We introduce a benchmark called Nullsettes, which assesses how well models can predict in silico loss-of-function (LOF) mutations, in synthetic expression cassettes with little evolutionary precedent.
+Testing 12 state-of-the-art gLMs, we find that most fail to consistently detect these strong LOF mutations.
+All models show a sharp drop in predictive accuracy as the likelihood assigned to the original (nonmutant) sequence decreases, suggesting that gLMs rely heavily on pattern-matching to their evolutionary prior rather than on any mechanistic understanding of gene expression.
+Our findings highlight fundamental limitations in how gLMs generalize to engineered, non-natural sequences, and underscore the need for benchmarks and modeling strategies that prioritize functional understanding.
+Keywords: genomic language model, synthetic biology, mutation prediction, gene expression, regulatory genomics
+1 Background
+Genomic language models (gLMs) learn a probability distribution over naturally occurring DNA sequences, approximating the evolutionary plausibility of a sequence or its probability of arising and persisting under evolutionary constraints [1, 2].
+This plausibility reflects factors such as the presence of common regulatory motifs or similarity to known natural genomes.
+For example, gLMs tend to assign higher likelihoods to promoters with frequent natural motifs than to random synthetic promoters.
+Evolutionary plausibility can be a useful proxy for biological function, as gLMs have already demonstrated strong zero-shot performance in predicting mutation effects to support sequence design.
 In both genomic and protein language models (gLMs and pLMs), higher model likelihoods often correlate with improved functional properties.
-For example, pLM likelihoods can guide the design of higher affinity antibodies (Hie et al., 2024), more efficient base editors (He et al., 2024), and brighter fluorescent proteins (Zhang et al., 2025).
-Similarly, gLM likelihoods enable genome-wide prediction of variant effects (Benegas et al., 2023) and 5’ untranslated region (UTR) optimization (Chu et al., 2024).
-However, current evaluations for gLMs on mutant effect prediction focus primarily on natural sequences.
-In contrast, synthetic biology often requires the design of functional sequences with little or no evolutionary precedent, such as sequences that confer novel functions, avoid crosstalk with native machinery, or push expression levels beyond natural limits.
-Examples include ultra-strong synthetic promoters (Schlabach et al., 2010), miRNA-based regulatory circuits that achieve dosage-compensated gene expression using elements orthogonal to native miRNAs (Du et al., 2024), and engineered metabolic pathways for production of small molecule drugs (Yan et al., 2023).
-For these applications, it is unclear whether gLMs trained on natural genomes can generalize to synthetic constructs, especially in the absence of deep mutational scanning data for DNA.
-If successful, gLMs could support more systematic and scalable approaches to genetic design.
+For example, pLM likelihoods can guide the design of higher antibodies [3], more efficient base editors [4], brighter fluorescent proteins [5], and deep mutational scaning (DMS) tasks [6, 7].
+Similarly, gLM likelihoods enable genome-wide prediction of variant effects [8] and 5 untranslated region (UTR) optimization [9].
+However, it remains unclear whether gLMs trained solely on natural genomes understand DNA function well enough to generalize to synthetic constructs with little to no evolutionary precedent.
+Current evaluations for gLMs on mutation effect prediction focus primarily on natural sequences, such as variants within endogenous cis-regulatory elements and splice sites [10].
+In contrast, synthetic biology often requires the design of functional sequences with little evolutionary history, to solve industrial challenges, avoid crosstalk with native pathways, or push expression levels beyond natural limits.
+Examples include ultra-strong synthetic promoters [11], miRNA-based regulatory circuits that achieve dosage-compensated gene expression using elements orthogonal to native miRNAs [12], and engineered metabolic pathways for production of small molecule drugs [13].
+We need evaluation benchmarks that test whether models can generalize beyond evolutionary patterns to capture functional principles to help rewire biology.
+Fundamental mechanisms of gene expression are conserved across all domains of life, making it an ideal test of functional understanding.
+We introduce, Nullsettes, a benchmark for evaluating whether gLMs can predict loss-of-function (LOF) mutations in synthetic expression cassettes.
+We create these LOF mutations in silico by rearranging key regulatory elements, such as promoters and start codons, within nonmutant cassettes.
+These rearrangements disrupt the canonical sequence structure required for transcription and translation, rendering the sequences non-functional.
+We curate nonmutant cassettes from massive parallel reporter assay (MPRA) datasets  [14, 15, 16, 17].
+GLMs assign low likelihood to these cassette as they contain genetic elements from evolutionarily distant species, and incorporate random sequences functioning as promoters.
+Across 12 state-of-the-art models (and 28 variants) spanning a wide range of design choices, we find that gLMs perform poorly on zero-shot LOF prediction, particularly under contexts that deviate strongly from natural sequence statistics.
+For example, model performance worsens when mutations cause more severe functional disruption, or when cassettes include strong promoters drawn from random sequence libraries.
+Notably, all models show a sharp drop in prediction accuracy as the likelihood of the nonmutant sequence decreases, indicating that gLMs rely heavily on pattern-matching to their evolutionary priors rather than on functional understanding of gene expression.
+Accurate predictions only emerge when the nonmutant likelihood surpasses a length-scaled threshold.
+Comparisons between models show that functional generalization may depend more on the relevance and quality of pretraining data than on the sheer size of the dataset.
+These findings highlight fundamental limitations in how current gLMs generalize to engineered sequences, and underscore the need for benchmarks and modeling strategies that prioritize the functional understanding of DNA sequences.
+2 Results and discussion
 Figure 1: Overview of Nullsettes.
 A) Schematic examples showing transcription (i) and translation (ii) becoming impaired when the order of control element such as terminator (iii) or stop codon (iv) are altered.
-B) Nullsettes generates 11 and 19 mutant variants for each eukaryotic and prokaryotic expression cassette, respectively, by translocating control elements, and evaluates how well gLMs can predict these loss-of-function mutations based on changes in log-likelihood.
-We introduce a benchmark suite, Nullsettes, to evaluate gLMs on their ability to predict loss-of-function (LOF) mutations in synthetic expression cassettes.
-Nullsettes leverages well-established mechanisms of gene expression to systematically introduce virtual LOF mutations by rearranging key control elements, such as promoters and start codons, within functional cassettes.
-Using functional cassettes (referred to as nonmutant) curated from Massive Parallel Reporter Assay (MPRA) data  (Lagator et al., 2022; de Boer et al., 2020; Kosuri et al., 2013; Zahm et al., 2024), including those with randomly generated, low-likelihood promoters, we show that Evo-2-7B outperforms 11 other state-of-the-art models in mutant effect prediction.
-However, all gLMs exhibit a sharp decline in predictive accuracy as the likelihood of the nonmutant sequence decreases.
-Moreover, the likelihood range at which each model performs optimally varies strongly with sequence length.
-These results suggest that both likelihood and length must be considered when applying gLMs to guide genetic design.
+B) Nullsettes consist of 11 and 19 mutant variants for each eukaryotic and prokaryotic expression cassette, respectively, by translocating control elements, and evaluates how well gLMs can predict these loss-of-function mutations based on changes in log-likelihood.
+We curate functional nonmutant cassettes from five MPRA datasets.
+C) Cumulative success rates of representative gLMs evaluated on five MPRA datasets, with promoter either being random sequences or based on natural promoter (indicated in bracket).
+The two yeast datasets use promoter containing both natural motifs (Abf1-TATA, polyT-polyA) and random sequences.
+Dashed lines denote 100% success on one dataset.
+Models are sorted by cumulative success rate.
+D) Success rates of models in identifying Nullsettes mutants for E.
+coli expression cassettes with either naturally-derived or random promoters.
+Each line connects the performance of a single model across the two sets of cassettes.
+Statistical significance with $p<0.05$ for a one-sided paired T-test.
+E) Scatter plot depicting the relationship between the number of transcriptional and translational steps affected (x-axis) and the number of models that successfully assign low pseudo-likelihood (y-axis) across Nullsettes mutants.
+Each point represents a distinct type of Nullsettes mutant, with eukaryotic (n = 11) and prokaryotic (n = 19) variants shown in blue and red, respectively.
+Dashed lines indicate linear regression fits for each group.
+We design Nullsettes as a benchmark to test whether genomic language models (gLMs) understand functional gene expression in sequences that lack evolutionary plausibility.
+Nullsettes consists of zero-shot mutation effect prediction tasks, where synthetic expression cassettes are subjected to virtual loss-of-function (LOF) mutations created by rearranging key genetic elements.
+In a canonical eukaryotic expression cassette (Figure 1A, i–ii), control elements including the promoter, start codon, coding sequence (CDS), stop codon, and terminator must appear in a specific 5-3 order to enable proper transcription and translation.
+Rearranging these elements, for example, placing a terminator upstream of the CDS (iii) or positioning the CDS downstream of the stop codon (iv), disrupts this structure and results in nonfunctional gene expression.
+Nullsettes focuses on single-element translocations that eliminate functional protein expression and cannot be rescued by any upstream or downstream sequences (Figure 1B).
+From these criteria, we obtain 11 mutation types for eukaryotic and 19 for prokaryotic cassettes, with additional prokaryotic variants involving rearrangement of ribosome binding site (RBS) (Supplementary Table S1S2).
+To ensure these tasks test generalization beyond evolutionary prior, we generate virtual mutants from synthetic expression cassettes with low gLM-assigned likelihoods but strong gene expression in either E.
+coli, yeast, or human, drawn from five massively parallel reporter assay (MPRA) datasets [14, 15, 16, 17] (Figure 1B).
+These cassettes combine sequences from distantly related species, for example, GFP as CDS from Aequorea victoria with promoters and terminators from bacteria, human, or mouse.
+Many cassettes even use random (but functional) promoters, which further lower gLM-assigned likelihood (Supplementary Figure S1).
+Together, these cassettes enable evaluation of model generalization to functional yet evolutionarily implausible sequences across species.
+Given all cassettes in a dataset, we evaluate mutation effect prediction in a zero-shot setting by comparing the distribution of nonmutant log-likelihood (LL) to the distribution of corresponding Nullsettes mutants (Figure 1B).
+We consider a model capable of identifying a mutation type if the mutant LL are significantly lower than the nonmutant LL, as determined by a paired t-test.
+A model’s success rate on the dataset is the proportion of mutation types identified.
+Figure 2: Relationship between gLM likelihood and zero-shot performance on mutation effect prediction.
+A) Each point represents an expression cassette.
+Scatterplots show the Nullsettes prediction performance of Evo2-7B and NT-2.5B-MS for cassettes from five datasets.
+B) Heatmap showing correlation between the gLM log likelihood (LL) of a sequence and gLM success rate on predicting its corresponding Nullsettes, across five datasets and 12 models.
+C) Log-likelihood ($L$) threshold to achieve a success rate >0.85>0.85 across datasets of increasing sequence length.
+Boxplots show the distribution of $L$ thresholds for top performing models (GENERanno, Evo2-7B, and Evo1.5) across five datasets, rescaled by lowest threshold.
+Datasets are ordered according to their average sequence length.
+Red dashed line is linear regression fit to the median threshold for each dataset.
+Across five datasets, we found that most genomic language models perform poorly in predicting LOF mutants in Nullsettes.
+We evaluated 12 representative models (and 28 variants; Supplementary Table S3S4, and S5) spanning state-of-the-art approaches to genomic language modeling.
+Nine out of twelve models achieved $50\%$ success rate on at least one dataset, failing to identify strong LOF mutations (Figure 1C).
+Only Evo2-7B and GENERanno-0.5B deliver strong, consistent performance (Supplementary Figure S2).
+GENERanno matched Evo2-7B performance despite having 14-fold fewer parameters and using 12-fold less pretraining data, potentially due to its curated pretraining on actively expressed regions of the genome [18], suggesting that data quality and relevance may matter more than scale for functional generalization.
+Supporting this hypothesis, performance did not consistently improve with scaling across four model families (Supplemental Figure S3).
+Current gLMs recognize what natural expression cassettes look like, but not how they work, failing to generalize to contexts that deviate strongly from natural sequence statistics.
+Across most models, performance drops significantly when identifying mutants with strong promoters drawn from random sequence libraries compared to promoters derived from nature (p =0.04=0.04)(Figure 1D).
+When evolutionarily plausibility is not a useful prior for function, as these random sequences are assigned significantly lower LL (Supplementary Figure S1), models struggle to predict mutation effects.
+A similar failure emerges across Nullsette mutations, with model performance worsening as mutations disrupt more steps of the gene expression process (Figure 1E).
+If models understood how genetic elements mechanistically drive expression, their predictions should remain robust or even improve as mutations become more disruptive.
+Instead, performance declines, suggesting that models do not reason about functional mechanism, and behave sensibly only within the distribution of natural sequences.
+We quantify disruption by counting the transcribed and translated elements missing in each mutant.
+When only one step is affected (e.g., a premature stop codon), the mutant remains close to natural sequence space, and models can often detect the resulting loss of function.
+But when multiple steps are disrupted (e.g., loss of both transcription and translation), the sequence falls further outside the distribution seen during training, and model outputs become unreliable.
+Taken together, these results show that current gLMs do not understand regulatory logic, generalizing based on sequence similarity to evolutionary prior.
+Across all gLMs, zero-shot performance in mutation effect prediction declines sharply as the log-likelihood (LL) of the original (nonmutant) sequence decreases, reinforcing that models rely on their evolutionary priors rather than a mechanistic understanding of DNA function.
+If a model understood how mutations disrupt gene expression, its predictions should depend on what function the mutation alters, not on how similar the sequence looks to the model’s training corpus.
+In that case, LOF mutants would be correctly identified regardless of the LL of the nonmutant cassette.
+However, we do observe this dependency on LL.
+Figure 2A shows this relationship for Evo2-7B and NT-2.5B-MS.
+For both models, the ability to detect LOF mutations improves consistently as the LL of the nonmutant increases.
+For example, in the E.
+coli (random) dataset [14], NT-2.5B-MS fails almost completely when LL <27<-27, but performs well above $80\%$ when LL >22>-22.
+This dependency is not specific to a single model.
+As shown in Figure 2B, the same correlation between nonmutant LL and prediction accuracy holds across nearly all models, regardless of architecture or training corpus, indicating a shared reliance on sequence prior rather than reasoning about expression mechanisms.
+Among top-performing models, the threshold LL required for high prediction accuracy is consistent and scales linearly with sequence length.
+Figure 2A shows that the minimum LL threshold for accurate prediction can vary widely across datasets.
+For example, NT-2.5B-MS achieves a success rate of $0.85$ at a threshold LL of approximately 20–20 for Yeast (Abf1TATA) cassettes, but around 16–16 for Human (natural).
+For each model–dataset pair, we estimate the LL threshold corresponding to a success rate of $0.85$ using linear regression.
+Figure 2C reveals that these thresholds are consistent among top performing models (after rescaling by lowest threshold).
+Furthermore, this threshold scales linearly with sequence length, showing a simple linear model can predict an effective LL threshold for zero-shot mutation effect prediction.
+3 Conclusions
+In this study, we introduce Nullsettes, a benchmark for evaluating the ability of genomic language models (gLMs) to understand DNA function by predicting loss-of-function (LOF) mutations in synthetic expression cassettes.
+Our analysis of 12 state-of-the-art models reveals that most struggle to consistently identify these LOF mutations, often relying on surface-level similarity rather than true functional understanding of regulatory elements.
+Reinforcing this fact, we show that LOF prediction accuracy declines sharply when nonmutant sequences have low gLM-assigned log-likelihood, as is the case for synthetic sequences.
+Comparing performance between models, we find a key avenue for improving functional understanding may be prioritizing curated, functionally relevant pretraining data over brute-force scaling.
+Together, these findings expose critical limitations in current gLMs and highlights the need for benchmark and modeling strategies that prioritize functional understanding for cell engineering.
 4 Methods
-2.1 Nullsettes Construction
-To systematically evaluate gLMs’ ability for zero-shot functional prediction on expression cassettes, we proposed Nullsettes, a suite of virtual mutants that perturb transcription and translation of an expression cassette by shuffling the order of six key control elements.
-These elements are promoter, ribosome binding site (RBS), start codon, coding sequence (CDS), stop codon, and terminator.
-In the canonical configuration of an expression cassette (Figure 1A i-ii), the proper ordering of promoter–start codon–CDS–stop codon–terminator from 5’ to 3’ end enables proper transcription (i) and translation (ii).
-However, translocations that reposition critical elements, such as placing the terminator between the promoter and the CDS (iii) or the CDS downstream of the stop codon (iv), result in transcriptional or translational failure, respectively.
-Nullsettes consists of single-element translocations that effectively eliminates any production of functional mRNA and protein (Figure 1B).
-Among the set of all possible single-element translocations, we subset for mutants where 1) the CDS expression cannot be rescued by any flanking sequences outside the cassette, 2) expression machinery can still act to generate useless products.
-These virtual mutant cassettes collectively make up the Nullsettes, which consists of 11 mutations for eukaryotic and 19 mutations for prokaryotic cassettes (see Appendix Table 3 and 4 for full list).
-The 8 additional mutants for prokaryote cassettes are due to translocating the RBS.
-Given a functional cassette, we then compare its gLM log-likelihood (LL) with its corresponding Nullsettes to assess how well a gLM can perform zero-shot functional prediction by consistently predicting lower likelihoods for the Nullsettes.
 4.1 Functional cassette curation
-To assess mutant effect prediction using Nullsettes, we curate expression cassettes from public MPRA datasets  (Lagator et al., 2022; de Boer et al., 2020; Kosuri et al., 2013; Zahm et al., 2024).
-We specifically use cassettes with strong gene expression but low gLM likelihood representing low evolutionary plausibility.
-These cassettes have low likelihood due to containing elements from evolutionarily distant organisms.
-For example, the CDS, green fluorscent protein (GFP), is from the jellyfish Aequorea victoria, whereas promoter and terminators are sourced from bacteria, human and mice.
-Furthermore, cassettes curated from Lagator dataset and deBoer pTpA dataset use random DNA sequences as promoters, which further lowers gLM likelihood compared to those with natural promoters/motifs (See Appendix A.3).
-In the paper, we name datasets with random promoters with the suffix (“Low”) and those with more natural promoters with the suffix (“High”) (Figure 2).
-Figure 2: Relationship between gLM likelihood and zero-shot performance on Nullsettes prediction.
-A) Each point represents an expression cassette.
-Scatterplots show the Nullsettes prediction performance of Evo-2-7B and Nucleotide Tranformer-2.5B-MS for sequences from five datasets.
-B) Heatmap showing correlation between the gLM log likelihood (LL) of a sequence and gLM success rate on Nullsettes prediction, across five datasets and 12 models.
-C) Stacked bar plot showing success rate of models computed using all sequences in each dataset, representing the proportion of Nullsettes mutants that a model consistently outputs a lower LL compared to the nonmutant.
+4.1.1 Prokaryote
+To evaluate the regulatory grammar of prokaryotes on genomic language models, we utilized two benchmark datasets derived from synthetic expression constructs in Escherichia coli.
+The first, the Kosuri dataset [16], comprises combinatorial assemblies of promoters and ribosome binding sites (RBS) upstream of a superfolder GFP (sfGFP) reporter, designed to dissect transcriptional and translational contributions to gene expression.
+The second, the Lagator dataset [14], features a large-scale library of random promoter sequences with experimentally measured activity, enabling evaluation of models’ ability to generalize to highly diverse, out-of-distribution regulatory inputs.
+4.1.2 Eukaryote
+To evaluate the capacity of genomic language models to generalize to eukaryotic regulatory logic, we utilized two large-scale massively parallel reporter assay (MPRA) datasets.
+The Zahm dataset [17] comprises a library of 6,144 synthetic promoters constructed by combining transcriptional response elements (TREs) from 229 human and mouse transcription factors with minimal promoters.
+These constructs were assayed across multiple human cell lines and stimulatory conditions to quantify dynamic, stimulus-specific gene expression.
+The deBoer dataset [15] includes a comprehensive collection of 100 million randomized promoter sequences in yeast, enabling high-resolution dissection of cis-regulatory grammar under out-of-distribution scenario.
+4.1.3 Promoter selection
+To construct decoupled expression cassettes, we selected 1,500 promoters or promoter–RBS pairs from four distinct datasets.
+For the Kosuri dataset [16], we selected candidate promoter–RBS pairs with protein expression levels exceeding $\mu+1.5\sigma$ (10,165) across the full distribution.
+In contrast, given the limited number of highly active constructs in the Lagator dataset, we directly selected the top 1,500 promoter sequences ranked by protein output.
+We then compared the log-likelihood (LL) distributions of promoter–RBS groups using genomic language models including Evo1-8k, GENERator-3B, GPN, and NT-2.5B-MS.
+As the Lagator dataset [14] consists of randomly generated promoters, whereas the Kosuri library is more rationally designed, we prioritized Kosuri promoter–RBS pairs whose LL distributions surpassed those from Lagator (see Supplementary Figure S1A), yielding 1,500 expression cassettes from each dataset.
+deBoer dataset [15] contains two promoter libraries: Abf1TATA and pTpA.
+Abf1TATA is designed by embedding conserved transcription factor binding sites such as Abf1 and a canonical TATA box, mimicking features of natural yeast promoters.
+In addition, pTpA consists of synthetic promoters constructed with a poly-T–poly-A architecture.
+Poly(dA-dT) sequences are naturally occurring motifs that associate with the centers of nucleosome-free regions in yeast promoters and serving as a hybrid natural + randomized library.
+We selectively constructed 1,500 active promoters for each dataset .
+The Zahm dataset [17] comprises synthetic promoters built by coupling one of three minimal promoters—minCMV, minProm, or minTK—with diverse transcriptional response element (TRE) units.
+TRE units contain Transcription factor binding motifs (TFBMs) mostly derived from human genomes.
+To construct a representative set of mammalian expression cassettes, we selected the top 1,500 promoter-TRE combinations exhibiting the highest transcriptional activity.
+4.2 Nullsettes Virtual Mutant Construction
+To assess whether genomic language models can detect syntactic violations in gene regulatory architecture, we constructed Nullsettes—a benchmark of completely non-functional expression cassettes.
+These sequences were generated by systematic component translocations that disrupt the canonical transcription–translation structure while retaining all original components.
+We first defined the canonical regulatory architecture for both prokaryote and eukaryote:
+Prokaryotic systems (e.g., E. coli) contain six ordered elements: Promoter - RBS - Start codon - CDS - Stop codon - Terminator.
+Eukaryotic systems (e.g., yeast, mammalian) consist of five ordered elements, lacking a RBS as translation initiation is guided by the 5 cap and the Kozak consensus sequence rather than a defined RBS: Promoter - Start codon - CDS - Stop codon - Terminator.
+For each system, we generated all possible single-component translocation mutants, in which one element was relocated to a different position while preserving the identity and total count of elements.
+To determine whether a mutant retained functional syntax, we evaluated each sequence against a minimal set of biologically grounded ordering constraints.
+These constraints ensure that proper CDS expression cannot be rescued by any flanking sequences outside the cassette.
+Specifically, we consider all circular permutations of each ordering and retain a mutant as a Nullsette only if no circular permutation of the mutant satisfies all three of the following rules:
+Rule 1: Proper initiation ordering — Ensures that transcription starts upstream of translation initiation and that the CDS is translated from an upstream start codon.
+Eukaryotic:  Position(Promoter) << Position(Start codon) << Position(CDS)
+Prokaryotic:  Position(Promoter) << Position(RBS) << Position(Start codon) << Position(CDS)
+Rule 2: Valid translation termination — The stop codon must follow the CDS or precede the start codon in circular order.
+(i) Position(CDS) << Position(Stop codon)  or  (ii) Position(Stop codon) << Position(Start codon)
+Rule 3: Valid transcription termination — The terminator must lie downstream of the CDS or upstream of the promoter in circular form.
+(i) Position(CDS) << Position(Terminator)  or  (ii) Position(Terminator) << Position(Promoter)
+After applying these rules across all circular permutations, we identified 19 non-functional translocation mutants in the prokaryotic system and 11 in the eukaryotic system, all of which violated the defined minimal functional syntax.
+These mutants are synthetically designed to be completely non-functional, challenging models to detect biologically catastrophic yet syntactically subtle disruptions in regulatory element ordering.
+All mutant sequences are listed in Supplementary Table S1 (prokaryote) and Supplementary Table S2 (eukaryote).
 4.3 Baseline Models
 We benchmarked a diverse set of 12 self-supervised genomic foundation models that represent current state-of-the-art approaches to DNA language modeling.
-In general, the models can be categorized based on their tokenization schemes (fixed-length k-mers, single-nucleotide tokens, byte-pair encoding, and hybrid schemes), pretraining strategies (masked language modeling (Devlin et al., 2019) vs.
-autoregressive modeling (Brown et al., 2020)), training corpus diversity (human references, plant genomes, multispecies genomes, and large-scale metagenomic assemblies), and architectural paradigms (CNN-based models, transformer-based models (Consens et al., 2025), and emerging architectures like StripedHyena (Poli et al., 2023; Ku et al., 2025) and Mamba (Gu & Dao, 2023)).
-Detailed model specifications are concluded in Appendix Table 1 and 2.
+In general, the models can be categorized based on their tokenization schemes (fixed-length k-mers, single-nucleotide tokens, byte-pair encoding, and hybrid schemes), pretraining strategies (masked language modeling [19] vs.
+autoregressive modeling [20]), training corpus diversity (human references, plant genomes, multispecies genomes, and large-scale metagenomic assemblies), and architectural paradigms (CNN-based models, Transformer-based models [2], and emerging architectures like StripedHyena [21, 22] and Mamba [23]).
+Detailed model specifications are concluded in Supplementary Table S4 for casual language modeling-based models and S5 for mask language modeling-based models.
 4.4 Evaluation Metrics
 We computed the mean base-pair log-likelihood (LL) as the sequence-level LL score.
-To account for the distinct pretraining objectives of causal language models (CLMs) and masked language models (MLMs), we applied different LL computation strategies, as detailed in Appendix B.1.
-After obtaining the LL distributions for both the original cassette and its virtual mutants (Nullsettes), we performed a one-sided paired permutation test to assess whether the mutant cassettes exhibit a significant decrease in LL compared to the original sequence (Figure 1B).
-Additional methodological details are provided in Appendix B.2.
-3 Results
-We benchmarked these gLMs and their variants across five datasets, as shown in Appendix C. Representative models were selected for further analysis in the following sections.
-3.1 Zero-shot Nullsettes prediction performance deteriorates for sequences with low gLM likelihood
-The evolutionary plausibility of a sequence, as predicted by a gLM, strongly correlates with the gLM’s performance in predicting mutation effects for the sequence.
-Figure 2A compares gLM prediction performance of individual functional cassettes and their predicted LL for two models, Evo-2-7B and NucleotideTransformer-2.5B-MS.
-Here success rate for each cassette is defined as the proportion of Nullsettes (11 in eukaryote, 19 in prokaryote) whose LL score is lower than the original, nonmutant cassette.
-For both models, we observe strong correlations between the nonmutant LL score and the gLM’s ability to identify LOF mutants.
-For example, for the E.coli (Low) dataset (Lagator et al., 2022), zero-shot Nullsettes prediction with NT-2.5B-MS is nearly impossible when the nonmutant LL is below -25, and easily accomplished for nonmutants with LL above -22.
-In fact, Figure 2B shows that all models, despite variations in architecture and training, exhibits a positive correlation between sequence likelihood and mutant prediction performance.
-The Evo-2-7B model achieves the best overall performance in zero-shot Nullsettes prediction.
-Evaluating zero-shot prediction using all sequences within a dataset (See Methods), Figure 2C shows that Evo-2-7B achieves the best overall performance across all five datasets.
-Specifically for the E.coli and yeast MPRA data using random promoters (left), which are less evolutionarily plausible than more natural sequences shown on the right (Appendix Figure 4).
-Figure 3: Optimal likelihood range for Nullsettes prediction varies with sequence length.
-A) Length in base pairs and normalized gLM log likelihood of all nonmutant cassettes, averaged across all 12 models.
-B) Contour of scatter plots similar to those in Figure 2A, colored by dataset, dotted vertical lines are rough estimates of optimal LL value for sequences from different datasets.
-Human data for Evo 1.5 is cut off for better visual due to large LL score.
-3.2 Optimal likelihood range predictive of strong model performance varies with sequence length
-Although mutant prediction performance correlates with sequence likelihood, the range of LL scores for which a model performs well depends highly on the length of the DNA sequence.
-The range of LL scores for which models achieve strong performance is higher for longer sequences, partly due to the fact that gLMs tend to assign higher LL scores to longer cassettes (Figure 3A).
-For example, Figure 3B shows that the NT-2.5B-MS model achieves strong performance for cassettes with log likelihood above $-22$ for both E.
-coli datasets, as indicated by the gray dotted line.
-However, for sequences from the Human GPRA dataset which are $2.5$-folds longer, only LL score above $-17$ achieved strong performance.
-Similarly sequences from the yeast (low) dataset are $30\%$ longer compared to sequences from the E.
-coli datasets.
-Thus, for GPN-promoter and Evo-1.5, the optimal likelihood range is also significant higher for the yeast (low) dataset.
-Altogether, there does not appear to be a single optimal likelihood range within which a model would be able to make zero-shot mutation effect predictions, as sequence properties such as length can have a significant effect on optimal likelihood range.
-4 Discussion
-In this study, we introduce Nullsettes, the first systematic benchmark for assessing the ability of genomic language models (gLMs) to predict loss-of-function mutations in synthetic expression cassettes.
-By enabling simple in silico translocation of sequences, Nullsettes facilitates generalizable, cassette-independent evaluations of gLM functional performance.
-Beyond general performance benchmarking, our results reveal a strong correlation between gLM performance on mutation prediction tasks and the model’s predicted likelihood of the original, non-mutant sequence.
-Moreover, we demonstrate that the optimal likelihood range for accurate mutant discrimination is not fixed but varies significantly with sequence length.
-Together, we hope this benchmark and its findings will inspire the development of gLMs that move beyond evolutionary priors to support the rational design of functional, out-of-distribution genetic sequences.
-Software and Data
-The source code for benchmarking gLM on Nullsettes is publicly available at: https://github.com/cellethology/GLM-Nullsette-Benchmark.
-Acknowledgements
-The authors thank members of the Cell Ethology Lab for valuable comments and suggestions.
-Impact Statement
-This paper presents work whose goal is to advance the field of genomic language model and its application in synthetic biology.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Datasets
-A.1 Prokaryote
-To evaluate the regulatory grammar of prokaryotes on genomic language models, we utilized two benchmark datasets derived from synthetic expression constructs in Escherichia coli.
-The first, the Kosuri dataset (Kosuri et al., 2013), comprises combinatorial assemblies of promoters and ribosome binding sites (RBS) upstream of a superfolder GFP (sfGFP) reporter, designed to dissect transcriptional and translational contributions to gene expression.
-The second, the Lagator dataset (Lagator et al., 2022), features a large-scale library of random promoter sequences with experimentally measured activity, enabling evaluation of models’ ability to generalize to highly diverse, out-of-distribution regulatory inputs.
-A.2 Eukaryote
-To evaluate the capacity of genomic language models to generalize to eukaryotic regulatory logic, we utilized two large-scale massively parallel reporter assay (MPRA) datasets.
-The Zahm dataset (Zahm et al., 2024) comprises a library of 6,144 synthetic promoters constructed by combining transcriptional response elements (TREs) from 229 human and mouse transcription factors with minimal promoters.
-These constructs were assayed across multiple human cell lines and stimulatory conditions to quantify dynamic, stimulus-specific gene expression.
-The deBoer dataset (de Boer et al., 2020) includes a comprehensive collection of 100 million randomized promoter sequences in yeast, enabling high-resolution dissection of cis-regulatory grammar under out-of-distribution scenario.
-A.3 Promoter selection
-To construct decoupled expression cassettes, we selected 1,500 promoters or promoter–RBS pairs from four distinct datasets.
-A.3.1 Kosuri and Lagator datasets
-For the Kosuri dataset (Kosuri et al., 2013), we selected candidate promoter–RBS pairs with protein expression levels exceeding $\mu+1.5\sigma$ (10,165) across the full distribution.
-In contrast, given the limited number of highly active constructs in the Lagator dataset, we directly selected the top 1,500 promoter sequences ranked by protein output.
-We then compared the log-likelihood (LL) distributions of promoter–RBS groups using genomic language models including Evo1 8k, GENERator 3B, GPN, and Nucleotide Transformer 2.5B multispecies.
-As the Lagator dataset (Lagator et al., 2022) consists of randomly generated promoters, whereas the Kosuri library is more rationally designed, we prioritized Kosuri promoter–RBS pairs whose LL distributions surpassed those from Lagator (Appendix Figure 4(a)), yielding 1,500 expression cassettes from each dataset.
-A.3.2 deBoer dataset
-deBoer dataset (de Boer et al., 2020) contains two promoter libraries: Abf1TATA and pTpA.
-Abf1TATA is designed by embedding conserved transcription factor binding sites such as Abf1 and a canonical TATA box, mimicking features of natural yeast promoters; while pTpA consists of synthetic promoters constructed with a simple poly-T–poly-A architecture, lacking specific transcription factor motifs and serving as a minimal, randomized control library.
-Similar to Kosuri vs Lagator, we selectively constructed 1,500 active promoters for each dataset based on LL distribution where pTpA is left-shifted comparing to Abf1TATA as Abf1TATA is more conservative (Appendix Figure 4(b)).
-The source can be found at NCBI’s GEO: GSE104878111https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE104878.
-A.3.3 Zahm dataset
-The Zahm dataset (Zahm et al., 2024) comprises synthetic promoters built by coupling one of three minimal promoters—minCMV, minProm, or minTK—with diverse transcriptional response element (TRE) units.
-To construct a representative set of mammalian expression cassettes, we selected the top 1,500 promoter-TRE combinations exhibiting the highest transcriptional activity.
-The source can be found at NCBI’s GEO: GSE271608222https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE271608.
-(a) Kosuri vs Lagator
-(b) deBoer TATA vs deBoer pTpA
-Figure 4: Comparison of promoter log-likelihood distribution
-A.4 Virtual mutant
-Translocation mutants were systematically engineered by permuting the canonical order of key regulatory and coding elements—promoter, ribosome binding site (RBS), start codon, coding sequence (CDS), stop codon, and terminator in prokaryotic systems; and promoter, start codon, CDS, stop codon, and terminator in eukaryotic systems.
-Each permutation aims to disrupt the natural transcription-translation flow, creating syntactically or functionally invalid constructs.
-A total of 19 translocation variants were designed for prokaryotic systems and 11 for eukaryotic systems.
-All constructs were rendered completely non-functional by design.
-All translocation combinations can be found in Appendix Table 3 for prokaryote and Appendix Table 4 for eukaryote.
-Appendix B Evaluation Metrics
-B.1 Mean Log-likelihood
-For casual language models (CLM), the log-likelihood of a nucleic acid sequence $X=(x_{1},x_{2},...,x_{n})$ is computed using logits output from the model.
+To account for the distinct pretraining objectives of causal language models and masked language models, we applied different LL computation strategies.
+Specifically, for casual language models (CLM), the log-likelihood of a nucleic acid sequence $X=(x_{1},x_{2},...,x_{n})$ is computed using logits output from the model.
 Since CLM operates in an autogressive manner, each token $x_{t}$ is predicted based on all previous token $x<t$.
@@
 Masked language models (MLMs) differ from autoregressive models in their inference behavior, as they are not inherently designed for sequential generation.
-To address this, Wang & Cho (2019) introduced the pseudo log-likelihood (PLL) approach, wherein each token in a sequence is masked individually to compute token-wise conditional probabilities.
+To address this, Wang and Cho [24] introduced the pseudo log-likelihood (PLL) approach, wherein each token in a sequence is masked individually to compute token-wise conditional probabilities.
 More recently, Gordon et al.
-(2024) proposed Single-Inference Pseudo Log-Likelihood, an efficient approximation that enables linear-time inference for BERT-style MLMs by exploiting training-time masking dynamics.
+[25] proposed Single-Inference Pseudo Log-Likelihood, an efficient approximation that enables linear-time inference for BERT-style MLMs by exploiting training-time masking dynamics.
 Rather than computing pseudo log-likelihood (PLL), we assessed relative changes in mean token-wise log-probabilities derived from unmasked logits for all MLM-based gLMs.
@@
 where all positions contribute to the likelihood computation since no masking is performed during inference.
-B.2 One-sided Paired Permutation Test
 To assess significance between paired conditions (e.g., original vs.
@@
 for the alternative hypothesis that the mutant is smaller than the original.
-Appendix C Model Performance Evaluation
-While representative models were selected for the main analysis, we conducted an extensive evaluation of various model versions across all datasets.
-As shown in Appendix Figure 5, the Evo series models (Nguyen et al., 2024; Merchant et al., 2024; Brixi et al., 2025) consistently outperformed all others in the Nullsettes tasks, with all the Evo variants ranking at the top.
-This advantage likely reflects the superior architectural design and scaling strategy employed by the Evo models.
-Among non-Evo models, METAGENE-1 (Liu et al., 2025b) achieved the highest performance, likely due to its training on over 1.5 trillion base pairs of DNA and RNA sequences derived from wastewater metagenomes.
-The unique breadth and diversity of metagenomic data may provide this model with an enhanced ability to generalize across sequence contexts.
-The Nucleotide Transformer (NT) (Dalla-Torre et al., 2024) series followed, ranking third overall.
-These findings highlight the impact of both model architecture and training data diversity on generalization performance across regulatory sequence prediction tasks.
-Figure 5: The number of cumulative successful predictions made by each model series across four datasets: Abf1TATA and pTpA (from deBoer), Zahm, Kosuri, and Lagator.
-Model name Pretraining dataset Pretraining method Tokenization Architecture Input length
-Nucleotide Transformer 2.5B multispecies (Dalla-Torre et al., 2024) 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e.
-roughly 29B tokens.
-MLM 6-mers Transformer, 2.5B params 1000 bp
-Nucleotide Transformer 500M multispecies (Dalla-Torre et al., 2024) 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e.
-roughly 29B tokens.
-MLM 6-mers Transformer, 500M params 1000 bp
-Nucleotide Transformer 500M Human ref (Dalla-Torre et al., 2024) GRCh38 human reference genome, resulting in 3B nucleotides, i.e. roughly 500M 6-mer tokens. MLM 6-mers Transformer, 500M params 1000 bp
-Evo1 8k (Nguyen et al., 2024) OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context
-Evo1 131k (Nguyen et al., 2024) OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens).
-CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 131,072 context using Evo1 8k as the base model
-Evo1.5 (Merchant et al., 2024) 50% increase in training data comparing to Evo1 CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context using Evo1 8k as the base model
-Evo2 1b base (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life.
-CLM Single-nucleotide tokens StripedHyena 2, 1B params Pretrained with 8192 context length
-Evo2 7b base (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life.
-CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 8192 context length
-Evo2 7b (Brixi et al., 2025) OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life.
-CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 1M context using Evo2 7b base as the base model
-GENERator 3b (Wu et al., 2025) Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate.
-CLM 6-mer Llama-based decoder, 3B params Pretrained with a context length of 98k bp
-GENERator 1.2b (Wu et al., 2025) Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate.
-CLM 6-mer Llama-based decoder, 1.2B params Pretrained with a context length of 98k bp
-METAGENE-1 (Liu et al., 2025b) 1.5T base pairs of DNA and RNA sequences from human wastewater samples. CLM Byte-pair encoding Llama-2, 7B params Pretrained with 512 sequence length
-Table 1: Genomic language model details
-Model name Pretraining dataset Pretraining method Tokenization Architecture Input length
-Caduceus-ph (Schiff et al., 2024) Human reference genome (35M tokens / nucleotide base pairs).
-MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (reverse complement augmentation)
-Caduceus-ps (Schiff et al., 2024) Human reference genome (35M tokens / nucleotide base pairs).
-MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (No reverse complement augmentation)
-GENA-LM human (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly. MLM Byte-pair encoding BERT/BigBird, 110M params About 4,500 nucleotides (512 BPE tokens)
-GENA-LM multi (Fishman et al., 2025) T2T vs.
-GRCh38.p13 human genome assembly augmented by sampling mutations from 1000-genome SNPs (gnomAD dataset) + Multispecies genomes from ENSEMBL release 108.
-MLM Byte-pair encoding BERT/BigBird, 110M params About 4,500 nucleotides (512 BPE tokens)
-GENA-LM large (Fishman et al., 2025) T2T vs. GRCh38.p13 human genome assembly. MLM Byte-pair encoding BERT/BigBird, 336M params About 4,500 nucleotides (512 BPE tokens)
-GPN (Benegas et al., 2023) Brassicales reference genome from NCBI Genome.
-Took the union of exons (with a small intronic flank), promoters (1,000 bp upstream of transcription start sites) as well as an equivalent amount of random windows from the whole genome.
-MLM Single-nucleotide tokens CNN 512 bp
-GPN-Promoter (Benegas et al., 2025a) Animal promoter, genomes of 434 animal species. MLM Single-nucleotide tokens ByteNet, 152M Pretrained on 512 bp sequences centered at TSSs of protein-coding genes
-DNABERT 2 (Zhou et al., 2023) Human genome dataset (2.75B nucleotide bases) + Multispecies genome dataset (from 135 species, spread across 6 cetogories).
-The dataset includes 32.49B nucleotides bases, excluding all sequence with N and retain only ATCG.
-MLM Byte-pair encoding BERT, 117M params Pretrained on 700 bp length sequences
-PDLLM-DNAMamba-6mer (Liu et al., 2025a) Plant reference genomes CLM 6-mer SSM, 130M params Pretrained max token length is 512
-Table 2: Genomic language model details (continued)
+5 Data availability
+The expression cassette of Kosuri dataset [16] was obtained from https://www.addgene.org/47441/.
+The Lagator dataset [14] was downloaded from https://github.com/szarma/Thermoters.
+The deBoer dataset [15] was downloaded from https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE104878.
+The Zahm dataset [17] was downloaded from https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE271608.
+6 Code availability
+The source code for benchmarking gLMs and the Nullsettes used in this work is publicly available at: https://github.com/cellethology/GLM-Nullsette-Benchmark.
+7 Acknowledgements
+The authors thank members of the Cell Ethology Lab for valuable comments and suggestions. This work was supported by the Westlake Fellows Program at Westlake University.
+8 Author Contributions
+Z.J.W.
+conceived the initial proposal.
+S.J.
+and Z.J.W.
+jointly developed the overall concept and contributed to the framework design.
+S.J.
+implemented the framework.
+S.J., X.L., and Z.J.W.
+conducted and analyzed the computational experiments.
+S.J.
+and Z.J.W.
+wrote the initial draft of the manuscript.
+9 Competing interests
+The authors declare no competing interests.
+Supplementary Information
+Supplemental Figures
+Figure S1: Comparison of promoter log-likelihood distribution
+Figure S2: The number of cumulative successful predictions made by each model series across four datasets: Abf1TATA and pTpA (from deBoer), Zahm, Kosuri, and Lagator.
+Figure S3: Number of Nullsette mutations successfully identified for gLMs with small and large variants, including Evo2 (1B vs.
+7B), GENERator (1.2B vs.
+3B), NT (500M vs.
+2.5B, multispecies), and gLM2 (150M vs.
+650M).
+Supplemental Tables
 Mutant ID Description
@@
 Translocation-19 Start codon - Promoter - RBS - CDS - Stop codon - Terminator
-Table 3: Prokaryotic virtual mutant cases. Translocation-3,4,5 cannot be compensated by a later start codon as there are no other in-frame start codon.
+Table S1: Prokaryotic virtual mutant cases. Translocation-3,4,5 cannot be compensated by a later start codon as there are no other in-frame start codon.
 Mutant ID Description
@@
 Translocation-11 Start codon - Promoter - CDS - Stop codon - Terminator
-Table 4: Eukaryotic virtual mutant cases. Translocation-2,3,4 cannot be compensated by a later start codon as there are no other in-frame start codon.
+Table S2: Eukaryotic virtual mutant cases. Translocation-2,3,4 cannot be compensated by a later start codon as there are no other in-frame start codon.
+Model series Representative variant
+METAGENE1 [26] METAGENE-1 (METAGENE1)
+Nucleotide Transformer [27] nucleotide-transformer-2.5b-multi-species (NT-2.5B-NT)
+GENERator [28] GENERator-eukaryote-3b-base (GENERator-3B)
+GENERanno [18] GENERanno-prokaryote-0.5b-base (GENERanno-0.5B)
+Evo1 [29, 30] evo-1.5-8k-base (Evo1.5)
+Evo2 [31] evo2_7b_base (Evo2-7B)
+DNABERT2 [32] DNABERT-2-117M (DNABERT2)
+Caduceus [33] caduceus-ph_seqlen-131k_d_model-256_n_layer-16 (Caduceus)
+GPN [8] gpn-brassicales (GPN)
+GPN-Promoter [34] gpn-animal-promoter (GPN-Promoter)
+gLM2 [35] gLM2_650M (gLM2)
+PDLLM [36] PDLLM-dnamamba-6mer (PDLLM)
+Table S3: Representative model variant.
+For each model series, we list a representative variant used in major benchmarking.
+The official model identifier is listed, followed in parentheses by the shorthand name used throughout the main text.
+For example, the variant “nucleotide-transformer-2.5b-multi-species” is referred to as “NT-2.5B-NT” in the text.
+Model name Pretraining dataset Pretraining method Tokenization Architecture Input length
+Evo1 8k [29] OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context
+Evo1 131k [29] OpenGenome: Prokaryotic whole-genomes dataset ( 300B tokens). CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 131,072 context using Evo1 8k as the base model
+Evo1.5 [30] 50% increase in training data comparing to Evo1 CLM Single-nucleotide tokens StripedHyena, 7B params Pretrained with 8,192 context using Evo1 8k as the base model
+Evo2 1B base [31] OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 1B params Pretrained with 8192 context length
+Evo2 7B base [31] OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life. CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 8192 context length
+Evo2 7B [31] OpenGenome2: a dataset containing 8.8 trillion tokens from all domains of life.
+CLM Single-nucleotide tokens StripedHyena 2, 7B params Pretrained with 1M context using Evo2 7b base as the base model
+GENERator eukaryote 3B [28] Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate.
+CLM 6-mer Llama-based decoder, 3B params Pretrained with a context length of 98k bp
+GENERator eukaryote 1.2B [28] Eukaryotic genomes ( 386B bp), plants, fungi, protozoa, mammlian, vertebrate, invertebrate.
+CLM 6-mer Llama-based decoder, 1.2B params Pretrained with a context length of 98k bp
+METAGENE-1 [26] 1.5T base pairs of DNA and RNA sequences from human wastewater samples. CLM Byte-pair encoding Llama-2, 7B params Pretrained with 512 sequence length
+PDLLM-DNAMamba [36] Plant reference genomes CLM 6-mer/5-mer/4-mer/3-mer/2-mer/single/BPE SSM, 130M params Pretrained max token length is 512
+Table S4: Casual language modeling (CLM) based genomic language models
+Model name Pretraining dataset Pretraining method Tokenization Architecture Input length
+Nucleotide Transformer 2.5B multispecies [27] 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e.
+roughly 29B tokens.
+MLM 6-mers Transformer, 2.5B params 1000 bp
+Nucleotide Transformer 500M multispecies [27] 850 whole genomes from NCBI, plants and viruses are not included, resulting in a total of 174B nucleotides, i.e.
+roughly 29B tokens.
+MLM 6-mers Transformer, 500M params 1000 bp
+Nucleotide Transformer 500M Human ref [27] GRCh38 human reference genome, resulting in 3B nucleotides, i.e. roughly 500M 6-mer tokens. MLM 6-mers Transformer, 500M params 1000 bp
+GENERanno prokaryote 0.5B base [18] 715 billion base pairs of prokaryotic DNA MLM single-nucleotide Transformer encoder, 500M params 8k bp
+GENERanno eukaryote 0.5B base [18] 386 billion base pairs of eukaryotic DNA MLM single-nucleotide Transformer encoder, 500M params 8k bp
+Caduceus-ph [33] Human reference genome (35M tokens / nucleotide base pairs).
+MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (reverse complement augmentation)
+Caduceus-ps [33] Human reference genome (35M tokens / nucleotide base pairs).
+MLM Single-nucleotide tokens BiMamba, 7.7M params Pretrained with 131,072 sequence length (No reverse complement augmentation)
+GPN [8] Brassicales reference genome from NCBI Genome.
+Took the union of exons (with a small intronic flank), promoters (1,000 bp upstream of transcription start sites) as well as an equivalent amount of random windows from the whole genome.
+MLM Single-nucleotide tokens CNN 512 bp
+GPN-Promoter [34] Animal promoter, genomes of 434 animal species. MLM Single-nucleotide tokens ByteNet, 152M Pretrained on 512 bp sequences centered at TSSs of protein-coding genes
+DNABERT2 [32] Human genome dataset (2.75B nucleotide bases) + Multispecies genome dataset (from 135 species, spread across 6 cetogories).
+The dataset includes 32.49B nucleotides bases, excluding all sequence with N and retain only ATCG.
+MLM Byte-pair encoding BERT, 117M params Pretrained on 700 bp length sequences
+gLM2 [35] OMG: encodes genomic scaffolds with both amino-acid (CDS) and DNA tokens (315B tokens).
+MLM Char-level tokens (DNA: lowercase, AA: uppercase) Transformer, 650M params / 150M params Pretrained with a 4096 token context window
+Table S5: Mask language modeling (MLM) based genomic language models